# Radiation forces and torques in optics and acoustics

### v1

@@ Abstract
Mechanical action of various kinds of waves have been known for several centuries. The first tide of scientific interest in wave-induced forces and torques emerged at the end of the 19th / beginning of the 20th centuries, with the development of wave theories and the concepts of wave momentum and angular momentum. A second tide appeared in the past several decades, connected to technological breakthroughs: the creation of lasers and the controlled generation of structured wavefields. This resulted in several important discoveries: optical trapping and manipulation of small particles, from atomic to micro sizes, as well as acoustic and acoustofluidic manipulation and sorting of larger particles, including biological cells and samples. Here we provide a unifying review of optical and acoustic forces and torques on various particles, addressing both their theoretical fundamentals and the main applications. Our approach employs the universal connection between the local energy, momentum, and spin densities in the wave fields and the principal forces and torques on small Rayleigh particles. Moreover, we describe the most important cases of nontrivial forces and complex particles: lateral and pulling forces, chiral and anisotropic particles, etc. We also describe the main experimental achievements and applications related to optical and acoustic forces and torques in structured wave fields. Our goal is to illuminate the common fundamental origin and close interconnections between the mechanical actions of optical and acoustic fields, in order to facilitate their profound understanding and the further development of optomechanical and acoustomechanical applications.
@@ I Introduction
The interaction between waves and matter is crucial in various areas of physics: from fluid mechanics to quantum field theory. The effect of matter on waves is obvious: for instance, a stone thrown to a water reservoir generates waves on the water’s surface. The reverse action of waves on matter is less obvious, but is also present for any kind of waves. For example, water-surface waves can shift bodies floating on the surface in the direction of wave propagation. Similar effects also appear for electromagnetic (light) and acoustic (sound) waves impinging on a material object.
Notably, for approximately 150 years, studies of optical and acoustic wave-induced mechanical effects have developed in parallel, progressing from 19th-century problems in radiation pressure and wave momentum, through advancements in optical and acoustic trapping and forces on small particles, to impressive modern achievements in optical/acoustic manipulation of particles, including ‘tweezers’, ‘spanners’, levitation, sorting, and 3D holographic arrangements.
In this review, we aim to provide a unified description of theoretical fundamentals and main experimental achievements in the mechanical action of electromagnetic and acoustic waves on various particles. We will show that these optical and acoustic effects have a common physical background, as well as numerous similar manifestations, and should be naturally considered within the same coherent approach.
@@ I.1 Historical overview
The idea of mechanical pressure exerted by light originates from Kepler’s observation in 1619 of a comet tail which was always directed away from the sun, as though blown by a ‘light-wind’ emanating from the sun [298], Fig. 1(a). In 1746, Euler wrote about a similar effect for sound waves in air: “sound vigorously excites not only a vibratory motion in the air particles, but one also observes a real motion in small, very light dust particles which tumble in the air” and immediately extended this idea to light: “it cannot be doubted that the vibratory motion caused by the light produces a similar effect” [572].
@@
According to Newtonian mechanics, any force on a mechanical body implies a momentum transfer, which leads us to think that light, or any other wave pushing the body, must carry momentum. The closely related concepts of light pressure and momentum were elaborated within Maxwell’s electromagnetic theory. In 1873, Maxwell wrote: “in a medium in which waves are propagated, there is a pressure in the direction normal to the waves, and numerically equal to the energy in unit of volume” [370]. In 1884, Poynting introduced the concept of electromagnetic energy flow (Poynting vector) [430], and in 1905 he described a beam of light as “a stream of momentum, the direction of the momentum being along the line of propagation, and the amount of momentum passing per second through unit area cross-section of the beam being equal to the density of the energy in it” [431, 432]. Remarkably, ten years before Poynting, in 1874, Umow described an analogous energy flow for acoustic waves in fluids and solids [552].
@@
Alongside with these theoretical developments, there were numerous attempts to measure radiation pressure in a laboratory. While the presence of a radiation pressure force from sound was already obvious to Euler, detection of light pressure was a challenging problem. In 1874, Crookes reported measurements of the “attraction and repulsion resulting from radiation” [159] using a ‘light-mill’. This device is now known as the Crookes radiometer [580, 115], Fig. 2(a), and it is readily available as an ‘executive toy’. It looks like a rotor inside a near-vacuum glass bulb, with its vanes having opposite white (reflecting) and black (absorbing) sides. When this rotor is exposed to light coming from one side, the white (black) vane surfaces reflect (absorb) light, which corresponds to double (single) amount of the momentum transfer, as shown in Fig. 2(a). This should produce a torque and rotation of the rotor proportional to the momentum of light. This rotation is indeed observed, but careful considerations show that it appears due to a thermal effect (photophoretic force) originating from heating of the black surfaces and motion of the heat-induced residual gas in the bulb [580, 115]. Furthermore, the resulting rotation often has opposite direction to that expected from the light-pressure-induced torque. Therefore, the Crookes radiometer failed to measure the radiation pressure. Almost 30 years later, in 1901, Lebedew, as well as Nichols and Hull, performed the first convincing laboratory measurements of electromagnetic radiation pressure consistent with Maxwell’s theory [323, 397, 398, 324].
@@
Remarkably, in 1877, near-simultaneously with the Crookes radiometer, Dvořák performed somewhat similar experiments with acoustic open (Helmholtz) resonators placed on a rotor and generating rotation when exposed to an external sound source [192], Fig. 2(b). This effect is readily reproduced in room conditions [275, 466]. It is usually explained by a flow of air induced by the sound and going out of the open side of the resonators (which is a nonlinear fluid-mechanical effect), but can also be described within linear acoustics and the asymmetric scattering diagram of the sound. The sound, scattered preferentially in the direction of the open end of the resonator, generates an oppositely-directed recoil force, as shown in Fig. 2(b). Note that an exact acoustic counterpart of the Crookes radiometer can also be constructed [171].
@@
In 1902–1905, almost simultaneously with the seminal works by Poynting, Rayleigh developed the theoretical fundamentals of the acoustic radiation pressure [446, 447], and envisaged the universal character of the wave-induced radiation forces: “it would be of interest to inquire whether other kinds of vibration exercise a pressure, and if possible to frame a general theory of the action” [446].
@@
Next, alongside the electromagnetic wave momentum and the corresponding radiation force, it was noticed that a circularly (or, generally, elliptically) polarized electromagnetic wave carries an intrinsic angular momentum (now known as spin) which can exert mechanical torque in interaction with bodies. The main features of these phenomena for light passing through an anisotropic crystal plate were described in the pioneering 1899 work by Sadowsky [468]. Since this work was published in an obscure Russian journal, it is worth translating and quoting its key conclusions: “A crystal plate, ground parallel to the principal dielectric axes, which transforms the normally incident linearly-polarized electromagnetic waves into elliptically-polarized ones, tends to rotate in the direction opposite to the rotation of electromagnetic forces (i.e., fields) generated by the plate. … These forces only tend to align the plate in a certain way but not to give it a continuous rotation.” and “The forces developed during the transmission of circularly-polarized electromagnetic waves through a crystal plate, except for the plate, tend to lead it into continuous rotation in the direction of rotation of the electromagnetic forces in the incident waves. … The largest torque is obtained for the plate”. This work fully anticipated optical torques on anisotropic plates, first measured 37 years later [80, 81, 265], and in detail, for small anisotropic particles, a century later [213]. In 1909, using mechanical analogies, Poynting explicitly suggested that circularly-polarized light carries angular momentum [433].
@@
However, despite these seminal historical works, optical and acoustic torques have not been investigated systematically until recently [111]. This is especially true for sound waves which cannot have circular polarizations in the paraxial plane-wave-like regime. Although the presence of intrinsic angular momentum (spin) produced by the acoustic counterparts of electromagnetic elliptical polarizations in structured (non-plane-wave) sound fields was anticipated in [290], the first experiment on the acoustic spin and torque was reported only in 2019 [488].
@@ I.2 Modern development
A revolution in the study of optical forces began with the invention of the laser in 1960 and the start of modern laser optics. In 1970, Ashkin reported the acceleration and trapping of micron-size dielectric particles in focused laser beams [36]. This discovery paved the avenue for the vast field of optical manipulation (trapping, tweezers, levitation, etc.) of particles which has an enormous range of applications. In particular, optical manipulation of atoms [439, 150, 427] and biological objects [41, 43] resulted in the Nobel Prizes in physics in 1997 and 2018. Notably, Ashkin’s pioneering work already mentioned the possibility of optical trapping of atoms and molecules, as well as “angular acceleration (i.e., torque) of trapped particles based on optical absorption of circularly polarized light”. Although he attributed the optical trapping to the “radiation pressure”, the main force responsible for it is the gradient force proportional to the gradient of the intensity of the electromagnetic field rather than the radiation-pressure force directed along the propagation (wave vector) of light. In fact, these two forces are near-orthogonal in paraxial optical beams.
@@
The gradient force was first described in 1957 by Gaponov and Miller for the case of charged particles trapped in a high-frequency electromagnetic field [222]. For neutral dipole (Rayleigh) particles, optical forces, including the gradient and radiation-pressure (scattering) parts, were considered in the 1970s by Gordon and Ashkin [243, 244, 44].
@@
Remarkably, small particles experience a very similar acoustic gradient force in a sound wave field, as was first described by Gor’kov in 1961 [245]. 20 years after Ashkin, Wu employed this force to achieve a stable acoustic trapping of latex particles and frog eggs in ultrasonic beams [583]. This started the rapidly developing field of acoustic manipulation with numerous applications, especially in life sciences and acoustofluidics [412, 375].
@@
Another revolution in optical and acoustic manipulation was provided by the development of structured optical and acoustic fields, which can have arbitrarily space-varying phase and amplitude profiles, essentially different from the simplest plane-wave approximation [248, 20, 463, 96]. An important example of such structured fields are vortex beams generated both in optics and acoustics in the 1990s [66, 16, 259]. Such beams have a doghnut-like transverse intensity profile and an azimuthal phase gradient producing an orbital angular momentum (OAM). Therefore, these beams are perfect for the radial trapping and azimuthal rotation of particles [258, 224, 161, 266, 50, 49].
Nowadays, structured optical and acoustic fields allow impressive 2D and 3D holographic manipulation of multiple particles [162, 53, 369, 373], Fig. 1(b,c), as well as efficient sorting of particles (e.g., biological molecules) based on their size, shape, and other parameters [412, 585, 318], Fig. 1(d). Recently, almost 150 years after the first experiments by Crookes and Dvorak, optical and acoustic forces trapping small particles were employed to produce volumetric displays, Fig. 3 [514, 263]. Clearly, optical and acoustic manipulations have much in common; yet, they nicely complement each other in their applicability ranges: nanometer to micron size particles for optical trapping and larger objects up to centimeter sizes for acoustic trapping [179]. For example, the optical and acoustic volumetric displays in Fig. 3 are based on the trapping of and particles, respectively.
@@
It should be noticed that most optical and acoustic manipulation research has been focused on radiation forces, while torques were explored almost marginally (apart from the OAM and vortex-beam context). Indeed, simple expressions for the dipole-approximation torques on isotropic Rayleigh particles in arbitrary structured fields were derived only recently [138, 126, 87, 401, 543]. Although radiation forces are indeed more important for trapping and tweezers, local torques have the same order of magnitude and can be crucial for rotations and 3D orientation of particles. Moreover, experiments with radiation torques on optically trapped particles in vacuum recently achieved GHz-scale rotation speeds [450, 8] reaching the quantum-oscillator regime [533, 241]. In turn, recent acoustic experiments with torques have produced kHz rotations of trapped particles [611].
@@
Let us recall that, from the very beginning, studies of optical and acoustic radiation forces (torques) were intimately related to the fundamental concepts of momentum (angular momentum) of the corresponding waves. For structured fields and waves in media, these concepts involved numerous controversies: the famous Abraham-Minkowski dilemma for the electromagnetic plane-wave momentum in a medium [426, 51, 380], the problem of the sound wave momentum [372, 418, 419], the separation of the spin and orbital angular momenta in optics [15, 21, 553] and field theory [322], etc.
@@
The past decade brought significant progress to these and related problems, particularly in the context of the mechanical action of structured wave fields. Namely, it was shown that in the case of monochromatic (single-frequency) and arbitrarily inhomogeneous optical and acoustic fields, one can define universal canonical momentum and spin angular momentum densities, such that the radiation pressure force and torque on isotropic Rayleigh particles are precisely determined by these densities [76, 87, 100, 321, 488, 543]. This provides a unifying and intuitively clear framework for the local wave momentum/force and spin/torque properties. Notably, the universality of this approach allows one to readily extend it to other types of waves, such as water-surface waves [103, 562].
@@
Importantly, the local spin density and the associated torque are inherent in this framework to any kind of classical waves: electromagnetic, sound, water, etc. This significantly extends the concept originally introduced by Sadowsky and Poynting, and for more than a century associated solely with electromagnetic waves. The explanation is as follows. While a plane electromagnetic wave can have a transverse circular polarization and spin, a plane sound wave involves purely longitudinal oscillations of the medium with no circular motion and spin. However, structured sound fields (i.e., superpositions of multiple plane waves with oscillations in different directions) generically produce elliptical local oscillations of the medium, similar to elliptical electromagnetic polarizations. This results in local spin angular momentum and its mechanical manifestations [290, 488, 101, 103].

### v3

@@ Abstract
The mechanical action of various kinds of waves has been recognized for several centuries. The first tide of scientific interest in wave-induced forces and torques emerged at the turn of the 20th century, with the development of wave theories and the concepts of wave momentum and angular momentum. A second surge occurred in the past several decades, driven by technological breakthroughs: the invention of lasers and the controlled generation of structured wave fields. This resulted in major discoveries, including optical trapping and manipulation of small particles, from atomic to micro sizes, as well as acoustic manipulation of larger particles, including biological cells and samples. Nowadays, radiation forces and torques underpin numerous applications: optical and acoustic tweezers, acoustofluidic sorting of biological cells, optomechanical systems operating in both classical and quantum regimes, solar sails, quantum simulators, volumetric displays, etc. In this review, we present a unifying perspective on optical and acoustic forces and torques acting on various particles, addressing both their theoretical foundations and key applications. Our approach relies on the universal connection between the local energy, momentum, and spin densities of wave fields and the principal forces and torques exerted on small particles. Moreover, we describe important cases of nontrivial (e.g., lateral and pulling) forces and complex (e.g., chiral and anisotropic) particles. We also highlight significant experimental achievements involving optical and acoustic manipulation in structured wave fields. Our aim is to illuminate the common fundamental origins and close interconnections between the mechanical actions of optical and acoustic fields, thereby fostering a deeper understanding and advancing the development of optomechanical and acoustomechanical applications.
@@ I Introduction
The interaction between waves and matter plays a fundamental role in many areas of physics, from fluid mechanics to quantum field theory. The effect of matter on waves is obvious: for instance, a stone thrown into a water reservoir generates waves on the water surface. The reverse action of waves on matter is less obvious, yet it also occurs for all types of waves. For example, water-surface waves can push floating bodies in the direction of wave propagation. Similar effects arise for electromagnetic (light) and acoustic (sound) waves impinging on material objects.
Notably, over the past 150 years, research on optical and acoustic wave-induced mechanical effects has developed in parallel. It has progressed from 19th-century studies of radiation pressure and wave momentum, through advances in optical and acoustic trapping and forces on small particles, to impressive modern achievements in particle manipulation, such as optical and acoustic ‘tweezers’, ‘spanners’, levitation, sorting, and 3D holographic arrangements.
In this review, we present a unified description of theoretical fundamentals and major experimental advances in the mechanical action of electromagnetic and acoustic waves on various particles. We demonstrate that these optical and acoustic effects share a common physical origin, exhibit many parallel manifestations, and can be naturally considered within the same coherent framework.
@@ I.1 Historical overview
The idea of mechanical pressure exerted by light dates back to 1619, when Kepler observed that a comet’s tail always points away from the Sun, as if blown by a ‘light wind’ emanating from it [317], Fig. 1(a). In 1746, Euler described a similar effect for sound waves in air: “sound vigorously excites not only a vibratory motion in the air particles, but one also observes a real motion in small, very light dust particles which tumble in the air”, and immediately extended this idea to light: “it cannot be doubted that the vibratory motion caused by the light produces a similar effect” [632].
@@
According to Newtonian mechanics, any force on a body originates from a momentum transfer, which leads us to think that light, or any other wave pushing the body, must carry momentum. The closely related concepts of light pressure and momentum were elaborated within Maxwell’s electromagnetic theory. In 1873, Maxwell wrote: “in a medium in which waves are propagated, there is a pressure in the direction normal to the waves, and numerically equal to the energy in unit of volume” [404]. In 1884, Poynting introduced the concept of electromagnetic energy flow (Poynting vector) [469], and in 1905 he described a beam of light as “a stream of momentum, the direction of the momentum being along the line of propagation, and the amount of momentum passing per second through unit area cross-section of the beam being equal to the density of the energy in it” [470, 471]. Remarkably, a decade earlier, in 1874, Umov described an analogous energy flow for acoustic waves in fluids and solids [611].
@@
Alongside these theoretical developments, there were numerous attempts to measure radiation pressure in a laboratory. While the presence of a radiation pressure force induced by sound was already obvious to Euler, detecting the pressure of light was far more challenging. In 1874, Crookes reported measurements of the “attraction and repulsion resulting from radiation” [168] using a so-called ‘light mill’ (now known as the Crookes radiometer) [640, 121], Fig. 2(a). This, now available as an ‘executive toy’, consists of a rotor inside a near-vacuum glass bulb, with its vanes painted white (reflective) on one side and black (absorptive) on the other. When this rotor is exposed to light coming from one side, the white (black) vane surfaces reflect (absorb) light, which corresponds to double (single) amount of the momentum transfer, as shown in Fig. 2(a). This should produce a torque and rotation of the rotor proportional to the light momentum. Rotation is indeed observed, but detailed analysis reveals that it arises from a thermal effect (photophoretic force) due to heating of the black surfaces and the resulting motion of the residual gas in the bulb [640, 121]. Furthermore, the observed rotation often occurs in the opposite direction to that expected from light pressure. Thus, the Crookes radiometer failed to measure the radiation pressure. Nearly 30 years later, in 1901, Lebedew, as well as Nichols and Hull, succeeded in making the first conclusive laboratory measurements of electromagnetic radiation pressure consistent with Maxwell’s theory [344, 434, 435, 345].
@@
Remarkably, in 1877, around the same time as the Crookes radiometer, Dvořák performed somewhat similar experiments using acoustic open (Helmholtz) resonators mounted on a rotor and generating rotation when exposed to an external sound source [202], Fig. 2(b). This effect is readily reproduced in room conditions [292, 512, 514]. It is usually explained by a flow of air induced by the sound and going out of the open side of the resonators (which is a nonlinear fluid-mechanical effect), but can also be described within linear acoustics and the asymmetric scattering of the sound. Namely, the sound, scattered preferentially in the direction of the open end of the resonator, generates an oppositely-directed recoil force, as shown in Fig. 2(b). Note that a direct acoustic analogue of the Crookes radiometer can also be constructed [180, 513].
@@
In 1902–1905, almost simultaneously with Poynting’s seminal works, Rayleigh developed the theoretical foundations of acoustic radiation pressure [486, 487]. He also envisaged the universal character of wave-induced radiation forces: “it would be of interest to inquire whether other kinds of vibration exercise a pressure, and if possible to frame a general theory of the action” [486].
@@
Next, alongside the electromagnetic wave momentum and the corresponding radiation force, it was noticed that a circularly (or, generally, elliptically) polarized electromagnetic wave carries an intrinsic angular momentum (now known as spin) which can exert a mechanical torque on bodies. The main features of this phenomenon for light passing through an anisotropic crystal plate were described in the pioneering 1899 work by Sadowsky [516]. Since this work was published in an antique Russian journal, it is worth translating and quoting its key conclusions: “A crystal plate, ground parallel to the principal dielectric axes, which transforms the normally incident linearly-polarized electromagnetic waves into elliptically-polarized ones, tends to rotate in the direction opposite to the rotation of electromagnetic forces (i.e., fields) generated by the plate. … These forces only tend to align the plate in a certain way but not to give it a continuous rotation.” and “The forces developed during the transmission of circularly-polarized electromagnetic waves through a crystal plate, except for the plate, tend to lead it into continuous rotation in the direction of rotation of the electromagnetic forces in the incident waves. … The largest torque is obtained for the plate”. Sadowsky’s work fully anticipated the optical torques on anisotropic plates that were first measured 37 years later [83, 84, 280], and—at a much finer scale—on small anisotropic particles a century later [226]. In 1909, using mechanical analogies, Poynting explicitly proposed that circularly-polarized light carries angular momentum [472].
@@
However, despite these early milestones, systematic investigations of optical and acoustic torques emerged only recently [116]. This is especially true for sound waves, which cannot exhibit circular polarizations in the paraxial, plane-wave-like regime. Although the presence of intrinsic angular momentum (spin) produced by the acoustic counterparts of electromagnetic elliptical polarizations in structured (non-plane-wave) sound fields was anticipated in 1973 [307], the first experimental demonstration of acoustic spin and torque was reported only in 2019 [540].
@@ I.2 Modern development
A revolution in the study of optical forces began with the invention of the laser in 1960 and the start of modern laser optics. In 1970, Ashkin reported the acceleration and trapping of micron-sized dielectric particles in focused laser beams [39]. This discovery paved the way for the vast field of optical manipulation, including trapping, tweezers, levitation, and related techniques, which now has an enormous range of applications. In particular, optical manipulation of atoms [478, 158, 466] and biological objects [44, 46] ultimately led to Nobel Prizes in physics in 1997 and 2018. Notably, Ashkin’s pioneering work already mentioned the possibility of optical trapping of atoms and molecules, as well as “angular acceleration (i.e., torque) of trapped particles based on optical absorption of circularly polarized light”. Although Ashkin attributed optical trapping to “radiation pressure”, the dominant mechanism is the gradient force proportional to the spatial gradient of the electromagnetic field intensity rather than the radiation-pressure force directed along the light’s propagation (wave vector). In fact, these two forces are near-orthogonal in paraxial optical beams.
@@
The gradient force was first described in 1957 by Gaponov and Miller for the trapping of charged particles in a high-frequency electromagnetic field [235]. For neutral dipole (Rayleigh) particles, optical forces, comprising both the gradient and radiation-pressure (scattering) parts, were examined in the 1970s by Gordon and Ashkin [256, 257, 47].
@@
Remarkably, small particles in a sound-wave field experience a very similar acoustic gradient force, first described by Gor’kov in 1961 [258]. Two decades after Ashkin, Wu employed this force to achieve a stable acoustic trapping of latex particles and frog eggs in ultrasonic beams [643]. This initiated the rapidly developing field of acoustic manipulation, with numerous applications, particularly in life sciences and acoustofluidics [450, 410].
@@
Another revolution in optical and acoustic manipulation came with the development of structured optical and acoustic fields, which can have arbitrarily space-varying phase and amplitude profiles, essentially different from the simplest plane-wave approximation [261, 449, 663, 22, 508, 100]. A prominent example is provided by vortex beams generated in both optics and acoustics in the 1990s [69, 17, 273]. Such beams feature a doughnut-shaped transverse intensity profile and an azimuthal phase gradient that produces orbital angular momentum (OAM). Vortex beams are perfectly suitable for radial trapping and azimuthal rotation of particles [272, 237, 170, 281, 53, 52].
Today, structured optical and acoustic fields enable impressive 2D and 3D holographic manipulation of multiple particles [171, 56, 401, 408], Fig. 1(b,c), as well as efficient sorting of particles (e.g., biological molecules) according to size, shape, and other parameters [450, 646, 339], Fig. 1(d). Recently, almost 150 years after the first experiments of Crookes and Dvorak, optical and acoustic forces have been employed to trap 3D arrays of small particles for volumetric displays, Fig. 3 [568, 278]. Furthermore, arrays of optically trapped atoms have provided a platform for quantum simulators [55, 56, 78, 203]. In addition, optical forces and torques underpin numerous optomechanical systems, spanning both classical and quantum regimes [48, 254, 413, 142, 588].
While optical and acoustic manipulations share many similarities, they also complement each other in their operational ranges: optical trapping is typically suited for particles from nm to 10 scales, whereas acoustic trapping excels with larger objects from 1 to 1 cm scales [188]. For example, the optical and acoustic volumetric displays in Fig. 3 are based on trapping particles of and , respectively.
@@
It should be noted that most research on optical and acoustic manipulation has focused on radiation forces, while torques have received comparatively little attention (apart from studies in the context of OAM and vortex beams). Indeed, simple expressions for the dipole-approximation torques on isotropic Rayleigh particles in arbitrary structured fields were derived only recently [145, 133, 91, 438, 602]. Although radiation forces are generally more important for trapping and tweezers, local torques have the same order of magnitude and can be crucial for inducing rotations and controlling orientation of the particles. Remarkably, recent experiments have achieved GHz-scale rotation speeds of optically trapped particles in vacuum [491, 9], reaching the quantum-oscillator regime [588, 254]. Likewise, recent acoustic experiments have produced kHz-scale rotations of trapped particles [676].
@@
Let us recall that, from the outset, studies of optical and acoustic radiation forces and torques have been closely tied to the fundamental concepts of wave momentum and angular momentum. For structured fields and waves in media, these concepts have sparked long-standing debates: the famous Abraham-Minkowski dilemma for the electromagnetic plane-wave momentum in a medium [465, 54, 415], the problem of the sound-wave momentum [406, 457, 458], and the separation of the spin and orbital angular momenta in optics [16, 23, 612] and in field theory [343], among others.
@@
The past decade has brought significant progress in resolving these issues, particularly in the context of the mechanical action of structured wave fields. It has been shown that for monochromatic (single-frequency) and arbitrarily inhomogeneous optical and acoustic fields, one can define universal canonical momentum and spin angular momentum densities [87], such that the radiation pressure force and torque on isotropic Rayleigh particles are precisely determined by these densities [79, 91, 105, 342, 540, 602]. This provides a unifying and intuitively clear framework for the local wave momentum/force and spin/torque properties. Notably, the universality of this approach makes it readily extendable to other types of waves, such as water-surface waves [108, 621].
@@
Importantly, in this framework, the local spin density and the associated torque are inherent to any kind of classical waves: electromagnetic, sound, water-surface, and others [87]. This significantly extends the concept originally introduced by Sadowsky and Poynting, which for many decades was associated exclusively with electromagnetic waves. The explanation is as follows. A plane electromagnetic wave can have a transverse circular polarization and the corresponding longitudinal spin angular momentum, whereas a plane sound wave consists of purely longitudinal oscillations of the medium, with no circular motion and spin. However, structured sound fields (i.e., superpositions of multiple plane waves with oscillations in different directions) generically produce elliptical local oscillations of the medium, similar to elliptical polarizations in electromagnetism. This results in local spin angular momentum and its mechanical manifestations [307, 540, 106, 108, 87].